OpenStudy (anonymous):

solve for the equation radical 7-radical x+3=2

OpenStudy (anonymous):

x+3=25

OpenStudy (anonymous):

x=22

OpenStudy (anonymous):

both of those answers are not correct

OpenStudy (anonymous):

\[\sqrt{7} - \sqrt{x+3} = 2\] First I would add the radical with x in it to the other side and subtract 2 to the other side to get: \[\sqrt{7} - 2 = \sqrt{x+3}\] then square both sides to get \[(\sqrt{7} - 2)^2 = x + 3\] With gets you \[7 -4\sqrt{7} + 4 = x + 3\] Leave x on that side and subtract 3 and combine like terms to get: \[8 - 4\sqrt{7} = x\] Factored down thats:\[4(2 - \sqrt{7}) = x\]